OpenStudy (anonymous):

Amelia is planning the trajectory of her next flight using the altitude and distance from Paris, France. She has determined her function to be f(x) = -4x + 102. Based on the situation above, describe any necessary restrictions to the domain and range. Provide justification for the restrictions.

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

You should not go to 0 or below. and shouldn't go above 50,000 miles

OpenStudy (anonymous):

i dont know how to solve this..

OpenStudy (anonymous):

What I told you is about what i know. Sorry

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

If you graph f(x) = -4x + 102, you'll find it has the y intercept of (0,102) and an x-intercept of (25.5, 0) You can find this x-intercept by plugging in f(x) = 0 and solving for x f(x) = -4x + 102 0 = -4x + 102 -102 = -4x -102/(-4) = x 25.5 = x x = 25.5

jimthompson5910 (jim_thompson5910):

So here is a blank xy axis 

OpenStudy (anonymous):

thats the answer?

jimthompson5910 (jim_thompson5910):

Then here is the graph 

OpenStudy (anonymous):

i dont need the graph though right?

jimthompson5910 (jim_thompson5910):

since x and f(x) must be positive, we know that we're only focusing on the first quadrant

jimthompson5910 (jim_thompson5910):

no, but it helps to have a graph to see what's going on

OpenStudy (anonymous):

ok so was that the answer you gave to me before the graph?